Mountain High Ice Cream Company transferred $78,000 of accounts receivable to the Prudential Bank. The transfer was made with recourse. Prudential remits 90% of the factored amount to Mountain High and retains 10% to cover sales returns and allowances. When the bank collects the receivables, it will remit to Mountain High the retained amount (which Mountain estimates has a fair value of $6,800). Mountain High anticipates a $4,800 recourse obligation. The bank charges a 2% fee (2% of $78,000), and requires that amount to be paid at the start of the factoring arrangement.

  

Required:

Prepare the journal entry to record the transfer on the books of Mountain High assuming that the sale criteria are met.
Answer #1

DR Cash $68,640
DR Loss on Sale of Receivables $7,360
DR Recievable from Factor $6,800
CR Recourse liability $4,800
CR Accounts Receivable $78,000

calculaions ::

(2% of $78,000) = 1,560

Dr Cash = (78,000 x .90) - (78,000 x .02) = > 70,200 - 1,560 = $68,640


Dr Loss on Sale of Rec = (4,800 + 78,000) - (68,640 + 6,800) = $7,360